# Combining Observational Data and Language for Species Range Estimation

### v1

@@ 1 Introduction
Species observation data, in the form of confirmed locations where the species have been observed to be present, have enabled biologists to create species range maps (SRMs) for a wide array of species. SRM provide valuable information for ecology, conservation, and environmental management. Recent advancements in generating SRMs have been achieved by combining vast amounts of observational data across thousands of species from platforms such as iNaturalist iNa and eBird Sullivan et al. [2009], and incorporating environmental covariates obtained from remote sensing data, climate models, geographic information systems. However, observational data can be difficult to gather due to geographic inaccessibility, resource constraints, and the lack of expertize required to identify species, particularly rare ones. Furthermore, rich environmental covariates might not be readily available across all geographic regions. As a result, current frameworks fall short in estimating SRMs for novel species, especially when observational data are scarce.
@@
We propose learning SRMs by combining observational data from citizen science observations with textual descriptions of species from Wikipedia (see Figure 1). These textural descriptions describe a wide array of properties including habitat preferences, range estimates, and visual attributes. Our framework learns to map location embeddings over the Earth’s surface and text embeddings from Wikipedia wik articles into a common space based on the species observations (see Figure 2). This provides a way to ground the information in text describing tens of thousands of species to spatial locations based on millions of observations. As shown in Figure 1, our resulting model allows us to estimate SRMs based on textual descriptions that might be known to an ecologist, such as habitat preferences, even when no location observations are available.
@@
We evaluate our text-driven approach for zero-shot estimation of SRMs for species from the IUCN IUC and eBird Status and Trends (S&T) Fink et al. [2020] benchmarks from Cole et al. [2023], which were excluded from the training data. Our model easily outperforms baselines (Table 1) and is competitive with SRMs estimated with ten observations for S&T species (Figure 3). Additionally, our model can be combined with observational data to achieve strong few-shot performance. Logistic regression models are regularised to be close to the species vector obtained from text descriptions, allowing us to match the performance of SRMs estimated using an order of magnitude more observations (Figure 3).
@@
While training is based on Wikipedia articles, it is unrealistic to expect a biologist to provide such extensive information for a novel species. Therefore, we evaluate our model based on short summaries, often just a few sentences long, that describe various aspects of the species. We further organize the summaries based on descriptions of the range (e.g., where they appear) and habitat preferences. While a biologist might not know the full range of an unknown species, they may be familiar with some of the latter (e.g., a tree might be in a tropical forest, or a bird might prefer wetlands). However, training on both forms of data allows the location embeddings to capture a wide range of geographic concepts embedded in language, such as countries, continents, climate regions, topology, and biomes, as revealed in our visualizations (Figures 4 and  5). Even short text summaries provide a significant boost in estimation accuracy over baselines when there are zero or few observations available.
@@ 3 Methods
Figure 2: LE-SINR learns to align location representations and text representations based on presence-only observation data and habitat or range descriptions for the corresponding species. Optionally, we can also include a learnable species embedding $E_{y}$ allowing range estimation for seen species $y$ from the training set. The model is trained on millions of observations from iNaturalist and language data from Wikipedia articles across thousands of species. LE-SINR supports zero-shot range estimation based on text descriptions for novel species and can also be used as a prior for few-shot range estimation.
@@ 3.1 Problem Setup
We focus on the problem of estimating SRMs across multiple species, indicating their presence or absence at each geographical location on Earth. Our observation dataset is composed of pairs $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}=(lat,lon)$ is a geographic location and $y_{i}\in\{1,\dots,S\}$ is the observed species ID. We use the dataset proposed in SINR Cole et al. [2023], consisting of 35.5 million observations covering 47,375 species observed prior to 2022 on the iNaturalist platform. Included species have at least 50 observations. Removing those that are included in the S&T and IUCN evaluation benchmarks results in 44,181 species, which forms our training set.
We also consider estimating SRMs from text-based descriptions in addition to observational data. To this end we curate a text dataset $\{(\mathbf{t}_{i},y_{i})\}$, where $\mathbf{t}_{i}$ are text descriptions sourced from Wikipedia articles of species $y_{i}$. For a particular species $y_{i}$ and corresponding text description $\mathbf{t}_{i}$, we aim to predict whether $y_{i}$ is present at some location $\mathbf{x}_{i}$ given $\mathbf{t}_{i}$. It is important to emphasize that this prediction only depends on the text description $\mathbf{t}_{i}$ and query location $\mathbf{x}_{i}$. Thus, we are not constrained to only making range predictions for species that have been observed during training. This allows for zero-shot species range estimation, where we can generate a range map of a species not in our observation dataset by utilizing a text description for it.
@@ 3.3 Language Enhanced SINR (LE-SINR) Architecture
We utilize the Spatial Implicit Neural Representation (SINR) Cole et al. [2023] framework for our approach, which models the probability of presence a species $y$ at given location $\mathbf{x}$ as $\sigma(f_{\theta}(\mathbf{x})\cdot E_{y})$, where $f_{\theta}(\mathbf{x})$ is a location encoding, $E_{y}$ is an embedding of species $y$ and $\sigma$ is the sigmoid function. Our architecture, shown in Figure 2, is similarly composed of two branches: one for representing locations and one for representing species. Each branch outputs a 256-dimensional feature vector. The probability of occurrence is then estimated by the sigmoid of their dot product.
The location branch is a location encoder model, $f_{\theta}(\mathbf{x})$, with parameters $\theta$, which takes a position embedding $\mathbf{x}$ (e.g., a location denoted by latitude and longitude) as input. The species branch is composed of two models, allowing us to generate species representations in two different ways. The first is a text-based species encoder, $g_{\phi}(\mathbf{t})$, with parameters $\phi$, which takes text $\mathbf{t}$ from our Wikipedia text data as input. The second species representation is a batch of tokens optimized directly, $E\in\mathbb{R}^{S\times 256}$. Given a known species $y\in\{1,\dots,S\}$ in the observation dataset, we can generate its representation with a simple lookup, $E_{y}$. Since we learn a unique token for each species in the training set, these species tokens cannot be used in the zero-shot setting. However, we are able to maintain the ability to have true supervised evaluation on species seen at training time. Additionally, the species tokens are used when a species has no text description.
Our text-based species encoder has two parts: a frozen LLM used to extract text embeddings and a learned fully connected network. As mentioned previously, to perform zero-shot SDM, we cannot directly utilize the ground truth species tokens. Instead, we input segments of text sourced from Wikipedia articles. Since these text segments can be as long as multiple paragraphs, we utilize a pretrained Large Language Model, GritLM Muennighoff et al. [2024], to create a fixed-length text embedding. GritLM is a recent language model with strong performance on text embedding tasks like document classification and retrieval. Due to the length of text in our data, it is much easier to work with a fixed-length embedding than with per-token embeddings from generative models like Llama3 lla . Given this text embedding, we then learn a three-layer fully connected language encoder network that outputs the final species embedding. As our approach incorporates language within the SINR framework for range estimation, we call it: Language Enhanced SINR (LE-SINR).
@@ 3.6 Implementation Details
Baselines Constant Prediction 0.01 0.22
Model Mean 0.09 0.35
Zero-shot results LE-SINR (Habitat Text) 0.28 0.51
LE-SINR (Range Text) 0.47 0.61
LE-SINR (Habitat Text) 0.32 0.52
LE-SINR (Range Text) 0.53 0.64
@@
Table 2: Ablation studies. Impact of various design choices, such as species description (range and habitat vs. taxonomy in LD-SDM Sastry et al. [2023]), location encoders, and the loss function on the zero-shot performance of LE-SINR. All models were trained without ‘Eval Sp’ and using ’Env’ when appropriate. LD-SDM is our implementation of Sastry et al. [2023].
LE-SINR Method IUCN S&T
LD-SDM (Genus Text)Sastry et al. [2023] 0.19 0.30
@@
GeoClip Backbone (Habitat text)Vivanco Cepeda et al. [2024] 0.23 0.49
@@
Spherical Harmonics Backbone (Habitat Text) Rußwurm et al. [2024] 0.31 0.52
Spherical Harmonics Backbone (Range Text) 0.53 0.63
SatClip Loss (Habitat Text) Klemmer et al. [2023] 0.26 0.52
@@
Proposed (Habitat Text) 0.32 0.52
Proposed (Range Text) 0.53 0.64
@@ 4.1 Zero-shot Range Estimation
The results show that the zero-shot range estimates comfortably outperform the baselines, achieving non-trivial performance with no observations. Range text works the best, but even habitat text achieves strong performance on both IUCN and S&T species. Similar to SINR, we observe a boost when including environmental covariates. LE-SINR with explicit species tokens matches the Oracle SINR performance, i.e.,, when evaluation species are used during training, suggesting that LE-SINR does not lose performance on observed species. Therefore, LE-SINR can be used both to explicitly model observed species and for zero-shot prediction from text for novel species.
@@
In Table 2 we compare against several other backbone architectures, loss functions, and training text. For the LD-SDM Sastry et al. [2023] baseline we use our LE-SINR architecture but replace the Wikipedia-derived text with LD-SDM taxonomic classification strings. For each species, we construct a string representing different levels of the taxonomic hierarchy, including class, order, family, genus, and species. During training, we randomly select one of these strings in each iteration. We then perform zero-shot evaluations at each taxonomic level. Performance decreases as we move up the hierarchy with species-level text performing the best. For the backbone baselines we replace our position encoder with the GeoClip Vivanco Cepeda et al. [2024] and Spherical Harmonics Rußwurm et al. [2024] architectures. We fine-tune the GeoClip model initialized with pre-trained weights. The spherical harmonics architecture is trained from scratch with $L=40$ polynomials. For the SatClip Klemmer et al. [2023] baseline, we use our model but replace the $AN_{\text{full}}$ loss with their contrastive loss applied to the position and text embeddings.
Figure 3: Range Estimation from Text and Observations. (Left) IUCN and (Middle) S&T results for zero-shot range estimation based on text, and few-shot estimation based on the text-driven prior. Both range and habitat texts improve few-shot performance over baseline SINR. (Right) Comparison of the position branch of SINR and LE-SINR for range estimation using a few examples. Language-driven covariates learned by LE-SINR lead to better generalization when observations are limited. We report the MAP for range estimation for species in the S&T and IUCN lists.
@@ 4.3 Evaluation of Learned Positional Embeddings
Prior work on generating SRMs often rely on carefully-selected spatial covariates such as temperature, elevation, precipitation, land cover, etc., to predict the geographical distribution of species based on a few observations. Our approach provides a way to learn a rich set of spatial covariates based on language data. Figure 3 (right) compares the position embeddings of SINR and LE-SINR as spatial covariates on the task of few-shot range estimation. In both cases, we use a model whose location branch is trained with position only (i.e., only latitude and longitude) as input to avoid conflation of the learned covariates with input ones. We find that the intermediate position embeddings learned when trained using language lead to better generalization, especially when training data is limited. Note that here we do not use any language input at test time, as both models are trained simply using observation data, same as the logistic regression baselines.
@@
Figure 5: Geospatial Grounding of Non-Species Concepts. Our model is able to geographically ground text prompts to locations on the earth. Here we display the inner product between the location encoder’s features and the language model’s encoding of the text displayed in the bottom left of each panel. This includes coarse concepts such as continents and countries (top row), geographic features such as specific lakes and mountain ranges (second row), in addition to concepts that do not appear in our species text training data but are likely already represented in the language model (third and fourth row). We do however observe some limitations resulting from the biases in our training data which favors North America, Europe, and Australasia (final row). Please zoom in to see more detail.
@@ 4.4 Limitations and Broader Impacts
Our models rely on text embeddings and summaries generated from LLMs, and thus may inherit the biases contained within them. For example, our model could further amplify biases in the data by incorrectly spatially localizing specific text terms inappropriately. Both Wikipedia text and observational data from iNaturalist are biased toward the United States and Western Europe. As a result, our models may not generalize as well to other geographical regions. As a result, there could be potential negative consequences associated with using the species range predictions from our model to inform conservation or policy decisions. While our results are promising, they may still fall short of the quality needed for such high-stakes use cases. Therefore, caution is encouraged when making decisions based on the model’s predictions.

### v2

@@ 1 Introduction
Collecting sufficient point-based observations of species in the wild allows us to infer species range maps (SRMs), which describe the spatial extent of where a species is likely to occur. These maps are invaluable, enhancing our understanding of natural history and informing land use and conservation decisions. Large-scale citizen science projects like iNaturalist iNa and eBird Sullivan et al. [2009] have recently accelerated SRM generation by systematically consolidating millions of observations across tens of thousands of species. By combining these extensive databases with environmental covariates, we can produce accurate SRMs. However, there remains a ’long tail’ of species with few observations, and current methods fall short of producing reliable range maps in such low-data settings.
@@
We propose learning SRMs by combining citizen science observations with text descriptions of species from Wikipedia (see Figure 1). These text descriptions describe a wide array of properties including habitat preferences, range estimates, and visual attributes. Our framework learns to map location embeddings over the Earth’s surface and text embeddings from Wikipedia wik articles into a common space based on the species observations (see Figure 2). This provides a way to ground the information in text describing tens of thousands of species to spatial locations based on millions of observations. As shown in Figure 1, our resulting model allows us to estimate SRMs based on text descriptions that might be known to an ecologist, such as habitat preferences, even when no location observations are available.
@@
We evaluate our text-driven approach for zero-shot estimation of SRMs for species from the IUCN IUC and eBird Status and Trends (S&T) Fink et al. [2020] benchmarks from Cole et al. [2023], which were excluded from the training data. Our model easily outperforms baselines (Table 2) and is competitive with SRMs estimated with ten observations for S&T species (Figure 3). Additionally, our model can be combined with observational data to achieve strong few-shot performance. Logistic regression models are regularized to be close to the species vector obtained from text descriptions, allowing us to match the performance of SRMs estimated using an order of magnitude more observations (Figure 3).
@@
While training is based on Wikipedia articles, it is unrealistic to expect a biologist to provide such extensive information for a novel species. Therefore, we evaluate our model based on short summaries, often just a few sentences long, that describe various aspects of the species. We further organize the summaries based on descriptions of the range (e.g., where they appear) and habitat preferences. While a biologist might not know the full range of an unknown species, they may be familiar with some of the latter (e.g., a tree might be in a tropical forest, or a bird might prefer wetlands). However, training on both forms of data allows the location embeddings to capture a wide range of geographic concepts embedded in language, such as countries, continents, climate regions, topology, and biomes, as revealed in our visualizations (Figures 4 and  5). Even short text summaries provide a significant boost in estimation accuracy over baselines when there are zero or few observations available. Our code and data is publicly available at: https://github.com/cvl-umass/le-sinr
@@ 3 Methods
Figure 2: LE-SINR learns to align location and text representations at training time using presence-only observation data and habitat or range descriptions for a set of species. Optionally, we can also include a learnable species token $E_{y}$ allowing range estimation for seen species $y$ from the training set. The model is trained on millions of observations from iNaturalist and language data from Wikipedia articles across thousands of species. LE-SINR supports zero-shot range estimation based on text descriptions for novel species and can also be used as a prior for few-shot range estimation.
@@ 3.1 Problem Setup
We focus on the problem of estimating SRMs across multiple species, indicating their presence or absence at each location on Earth. For a given location $\mathbf{x}$, our goal is to predict the probability of each species being present there. Our observation dataset is composed of pairs $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}=(lat,lon)$ is a geographic location and $y_{i}\in\{1,\dots,S\}$ is the observed species label. We use the dataset proposed in SINR Cole et al. [2023], consisting of 35.5 million observations covering 47,375 species observed prior to 2022 on the iNaturalist platform. Included species have at least 50 observations. Removing those that are included in the S&T and IUCN evaluation benchmarks results in 44,181 species, which forms our training set.
We also consider estimating SRMs from text-based descriptions in addition to observational data. To this end we curate a text dataset $\{(\mathbf{t}_{i},y_{i})\}$, where $\mathbf{t}_{i}$ is a text description sourced from Wikipedia pertaining to species $y_{i}$. For a particular species $y_{i}$ and corresponding text description $\mathbf{t}_{i}$, we aim to predict whether $y_{i}$ is present at some location $\mathbf{x}_{i}$ given $\mathbf{t}_{i}$. It is important to emphasize that this prediction only depends on the text description $\mathbf{t}_{i}$ and query location $\mathbf{x}_{i}$. Thus, we are not constrained to only making range predictions for species that have been observed during training. This allows for zero-shot species range estimation, where we can generate a range map of a species not in our observation dataset by utilizing a text description for it.
@@ 3.3 Language Enhanced SINR (LE-SINR) Architecture
We utilize the Spatial Implicit Neural Representation (SINR) Cole et al. [2023] framework for our approach, which models the probability of presence for a species $y$ at given location $\mathbf{x}$ as $\sigma(f_{\theta}(\mathbf{x})\cdot E_{y})$, where $f_{\theta}(\mathbf{x})$ is a location encoding, $E_{y}$ is an embedding of species $y$ and $\sigma$ is the sigmoid function. Our architecture, shown in Figure 2, is similarly composed of two branches: one for representing locations and one for representing species. Each branch outputs a 256-dimensional feature vector. The probability of occurrence is then estimated by the sigmoid of their dot product.
The location branch is a location encoder model, $f_{\theta}(\mathbf{x})$, with parameters $\theta$, which takes a position embedding $\mathbf{x}$ (e.g., a location denoted by latitude and longitude) as input. The species branch is composed of two models, allowing us to generate species embeddings in two different ways. The first is a text-based species encoder, $g_{\phi}(\mathbf{t})$, with parameters $\phi$, which takes text $\mathbf{t}$ from our Wikipedia text data as input. The second species representation is a batch of species tokens optimized directly, $E\in\mathbb{R}^{S\times 256}$. Given a known species $y\in\{1,\dots,S\}$ in the observation dataset, we can generate its representation with a simple lookup, $E_{y}$. Since we learn a unique species token for each species in the training set, these species tokens cannot be used in the zero-shot setting. However, we are able to maintain the ability to have true supervised evaluation on species seen at training time. Additionally, the species tokens are used when a species has no text description.
Our text-based species encoder has two parts: a frozen LLM used to extract text embeddings and a learned fully connected network. As mentioned previously, to perform zero-shot SDM, we cannot directly utilize the species tokens. Instead, we input segments of text sourced from Wikipedia articles. Since these text segments can be as long as multiple paragraphs, we utilize a pretrained Large Language Model, GritLM Muennighoff et al. [2024], to create a fixed-length text embedding. GritLM is a recent language model with strong performance on text embedding tasks like document classification and retrieval. Due to the length of text in our data, it is much easier to work with a fixed-length embedding than with per-token embeddings from generative models like Llama3 lla . Given this text embedding, we then learn a three-layer fully connected language encoder network that outputs the final species embedding. As our approach incorporates language within the SINR framework for range estimation, we call it: Language Enhanced SINR (LE-SINR).
@@ 3.6 Implementation Details
Baselines Constant Prediction 0.01 0.22
Model Mean 0.09 0.35
Zero-shot LE-SINR (Habitat Text) 0.28 0.51
LE-SINR (Range Text) 0.47 0.61
LE-SINR (Habitat Text) 0.32 0.52
LE-SINR (Range Text) 0.53 0.64
Table 2: LE-SINR Ablations. Impact of various design choices, such as species description (range and habitat vs. taxonomy in LD-SDM Sastry et al. [2023]), location encoders, and the loss function on the zero-shot performance of LE-SINR. All models were trained without ‘Eval Sp’ and using ’Env’ when appropriate. LD-SDM is our model and text encoder but trained with taxonomic text as in Sastry et al. [2023].
Ablation Type Method +Env IUCN S&T
Text LD-SDM Sastry et al. [2023] (Genus Text) 0.19 0.30
@@
Position Encoder GeoCLIP Backbone Vivanco Cepeda et al. [2024] (Habitat text) 0.23 0.49
@@
Spherical Harmonics Encoding Rußwurm et al. [2024] (Habitat Text) 0.31 0.52
Spherical Harmonics Encoding (Range Text) 0.53 0.63
Loss Function SatCLIP Loss Klemmer et al. [2023] (Habitat Text) 0.26 0.52
@@
Ours (Habitat Text) 0.32 0.52
Ours (Range Text) 0.53 0.64
@@ 4.1 Zero-shot Range Estimation
The results show that the zero-shot range estimates comfortably outperform the baselines, achieving non-trivial performance with no observations. Range text works the best, but even habitat text achieves strong performance on both IUCN and S&T species. Similar to SINR, we observe a boost when including environmental covariates. LE-SINR with explicit species tokens matches the Oracle SINR performance, i.e., when evaluation species are used during training, suggesting that LE-SINR does not lose performance on observed species. Therefore, LE-SINR can be used both to explicitly model observed species and for zero-shot prediction from text for novel species.
@@
In Table 2 we compare against other training text, position encodings, and loss functions. For the LD-SDM Sastry et al. [2023] baseline we use our LE-SINR architecture but replace the Wikipedia-derived text with LD-SDM taxonomic classification strings. Unlike LD-SDM’s autoregressive LLM, which generates an embedding for each token, this baseline, like our main method, encodes text using a model trained specifically for fixed-size text embeddings. For each species, we construct a string representing different levels of the taxonomic hierarchy, including class, order, family, genus, and species. During training, we randomly select one of these strings in each iteration. We then perform zero-shot evaluations at each taxonomic level. Performance decreases as we move up the hierarchy with species-level text performing the best. For the position encoder baselines we first replace our position encoder with the pre-trained GeoCLIP Vivanco Cepeda et al. [2024] model. The weights are then fine-tuned during training. We also try the Spherical Harmonics Rußwurm et al. [2024] coordinate encoding to encode the latitude and longitude. We experimented with multiple values for $L$ as well as the SIREN network architecture and found $L=10$ with our original MLP architecture worked best for this experiment. For the SatCLIP Klemmer et al. [2023] baseline, we use our model but replace the $AN_{\text{full}}$ loss with their contrastive loss applied to the position and text embeddings. Looking at Table 2, none of these experiments perform better than our method.
Figure 3: Range Estimation from Text and Observations. (Left) IUCN and (Middle) S&T results for zero-shot range estimation based on text, and few-shot estimation based on the text-driven prior. Both range and habitat texts improve few-shot performance over baseline SINR. (Right) Comparison of the position branch of SINR and LE-SINR for range estimation using a few examples. Language-driven covariates learned by LE-SINR lead to better generalization when observations are limited. We report the MAP for range estimation for species in the S&T and IUCN test sets.
@@ 4.3 Evaluation of Learned Positional Embeddings
Prior work on generating SRMs often rely on carefully-selected spatial covariates such as temperature, elevation, precipitation, land cover, etc., to predict the distribution of species based on a few observations. Our approach provides a way to learn a rich set of spatial covariates based on language data. Figure 3 (right) compares the position embeddings of SINR and LE-SINR as spatial covariates on the task of few-shot range estimation. In both cases, we use a model whose location branch is trained with position only (i.e., only latitude and longitude) as input to avoid conflation of the learned covariates with input ones. We find that the intermediate position embeddings learned when trained using language lead to better generalization, especially when training data is limited. Note that here we do not use any language input at test time, as both models are trained simply using observation data, same as the logistic regression baselines.
@@
Figure 5: Geospatial Grounding of Non-Species Concepts. LE-SINR is able to geographically ground text prompts to locations on the earth. Here we display the inner product between the location encoder’s features and the language model’s encoding of the text displayed in the bottom left of each panel. This includes coarse concepts such as continents and countries (top row), geographic features such as specific lakes and mountain ranges (second row), in addition to concepts that do not appear in our species text training data but are likely already represented in the language model (third and fourth row). We do, however, observe some limitations resulting from the biases in our training data which favors North America, Europe, and Australasia (final row). Please zoom in to see more detail.
@@ 4.4 Limitations and Broader Impacts
Another limitation is that our models are most applicable when range and habitat descriptions are available with very few observations, which is not common. While we evaluate our approach based on Wikipedia, a more realistic scenario involves obtaining descriptions of rarely observed species from domain experts. However, we note that many species have substantial Wikipedia articles despite having fewer than 50 observations on iNaturalist.
Our models rely on text embeddings and summaries generated from LLMs, and thus may inherit the biases contained within them. For example, our model could further amplify biases in the data by incorrectly spatially localizing specific text terms inappropriately. Both Wikipedia text and observational data from iNaturalist are biased toward the United States and Western Europe. As a result, our models may not generalize as well to other geographical regions. There could also be potential negative consequences associated with using the species range predictions from our model to inform conservation or policy decisions. While our results are promising, they may still fall short of the quality needed for such high-stakes use cases. Therefore, caution is encouraged when making decisions based on the model’s predictions. Another risk is that the model could be used to locate threatened or endangered species. To mitigate this, we only train on publicly available observation data that has been deemed safe to redistribute by the iNaturalist platform.